OpenStudy (anonymous):

Decide if the sequence is geometric. if yes, state the next 3 terms and the common ratio. 1) 6,24,96,384 ...... 2) 136,68,34,17, .... 3) 1/3, 2/3, 1 , 4/3 .......

OpenStudy (anonymous):

@jim_thompson5910 can you explain to me, im so confused on where to start

OpenStudy (anonymous):

the first one I believe is NO

jimthompson5910 (jim_thompson5910):

to figure out if it's a geometric sequence, you divide each term by the previous term

jimthompson5910 (jim_thompson5910):

if you get the same common ratio every time, then it's definitely geometric

jimthompson5910 (jim_thompson5910):

and once you figure out this common ratio, you use it to generate more terms

OpenStudy (anonymous):

to find more. I have to do 34 divide by 4 ? because the common ratio is 4. or how ?

jimthompson5910 (jim_thompson5910):

which one are you on?

jimthompson5910 (jim_thompson5910):

#2?

OpenStudy (anonymous):

number one. do I multiply 384 by 4 to get the next 3 terms ?

jimthompson5910 (jim_thompson5910):

yes you multiply 384 by 4 to get the next term then you multiply that term you just got by 4 to get the next term etc etc

jimthompson5910 (jim_thompson5910):

this of course only works if the sequence is geometric

OpenStudy (anonymous):

yes yes okay ! I got it .. what if its a fraction like numbe3 ?

jimthompson5910 (jim_thompson5910):

what is 2/3 divided by 1/3

OpenStudy (anonymous):

1/3?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

(2/3) divided by (1/3) = (2/3)*(3/1) = ??

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what is 1 divided by 2/3

OpenStudy (anonymous):

the same ?

jimthompson5910 (jim_thompson5910):

you sure?

OpenStudy (anonymous):

no ,

jimthompson5910 (jim_thompson5910):

1 divided by 2/3 = (1/1)*(3/2) = ??

OpenStudy (anonymous):

3/2?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so before you got 2, but now you're getting 3/2

jimthompson5910 (jim_thompson5910):

this means that the ratio has changed and there is no common ratio

OpenStudy (anonymous):

so its no

jimthompson5910 (jim_thompson5910):

yep, it's not a geometric sequence

OpenStudy (anonymous):

ok! I swear Im going to put you in my graduation speech because you've helped a lot .

jimthompson5910 (jim_thompson5910):

lol I'm flattered, but I haven't helped that much really

jimthompson5910 (jim_thompson5910):

glad it's all clicking though

OpenStudy (anonymous):

ha but anyways what if they ask for the 6th term?

jimthompson5910 (jim_thompson5910):

for which one

OpenStudy (anonymous):

4) 8,2, 1/2 common ratio I found is 4 .

jimthompson5910 (jim_thompson5910):

not 4, but close (sorta)

jimthompson5910 (jim_thompson5910):

2/8 = ???

OpenStudy (anonymous):

how you ge2/8

jimthompson5910 (jim_thompson5910):

i picked the second term and divided it by the previous term

OpenStudy (anonymous):

so you got 2/8 as the ratio?

jimthompson5910 (jim_thompson5910):

reduce it

jimthompson5910 (jim_thompson5910):

what do you get when you do

OpenStudy (anonymous):

1/4

jimthompson5910 (jim_thompson5910):

so what's what I meant when I said you were close

jimthompson5910 (jim_thompson5910):

the common ratio is 1/4 and not 4

jimthompson5910 (jim_thompson5910):

so to get the next term, you multiply the last term by 1/4

OpenStudy (anonymous):

I would multiply 1/2 and 1/4?

jimthompson5910 (jim_thompson5910):

multiply the numerators together multiply the denominators together

jimthompson5910 (jim_thompson5910):

1/2 times 1/4 = (1*1)/(4*2) = 1/8

OpenStudy (anonymous):

1/8

jimthompson5910 (jim_thompson5910):

yep correct

jimthompson5910 (jim_thompson5910):

1/8 times 1/4 = ???

OpenStudy (anonymous):

1/32

jimthompson5910 (jim_thompson5910):

and the term after 1/32 is ???

OpenStudy (anonymous):

128

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

what! it is 1/128

jimthompson5910 (jim_thompson5910):

yes, so that's the sixth term

OpenStudy (anonymous):

ok ! can you help on a word problem ?

jimthompson5910 (jim_thompson5910):

sure one more

OpenStudy (anonymous):

you got to go?

jimthompson5910 (jim_thompson5910):

no but I don't want to do all your hw although you seem far from the type that just dumps all their hw on here and doesn't bother to learn, so that's good

OpenStudy (anonymous):

oh yes I know. I just got two more and ill do the rest by myself.

jimthompson5910 (jim_thompson5910):

ok great, I don't want to discourage you from asking for help

jimthompson5910 (jim_thompson5910):

I just don't want you too dependent on this site either

OpenStudy (anonymous):

yes I understand. can I ask you something im always curious about lol

jimthompson5910 (jim_thompson5910):

sure go for it

OpenStudy (anonymous):

how old are you

jimthompson5910 (jim_thompson5910):

26 why do you ask

OpenStudy (anonymous):

just always been curious

jimthompson5910 (jim_thompson5910):

i gotcha

jimthompson5910 (jim_thompson5910):

I guess I don't sound like a teen huh?

jimthompson5910 (jim_thompson5910):

anyways, what's your other questions (math questions lol)

jimthompson5910 (jim_thompson5910):

then again, age is a number which is a math thing too, so that's partially relevant

OpenStudy (anonymous):

you just sound or seem, well educated and well mannered lol yes hold on its long

jimthompson5910 (jim_thompson5910):

hmm I guess so, but a teen could sound like it too maybe

jimthompson5910 (jim_thompson5910):

so it's a word problem?

OpenStudy (anonymous):

heavy rain in brieannes town caused the river to rise. the river rose three inches the first day. and each day after that it rose twice as much as the previous day. what type of sequence is own in the above situation? why is it this type of sequence? write a rule for the nthterm of this situation? how manydid the river rise the 7th day? howmuch total did the river rise in five days?

jimthompson5910 (jim_thompson5910):

so we basically have this sequence 3, 6, 12, 24, 48, 96, ... ie you start with 3 and you double each term to get the next term

jimthompson5910 (jim_thompson5910):

so what type of sequence is this?

OpenStudy (anonymous):

geometric

jimthompson5910 (jim_thompson5910):

how do you know this

OpenStudy (anonymous):

doubles each term

jimthompson5910 (jim_thompson5910):

so in more mathematical terms, you're basically saying the common ratio is r = 2

jimthompson5910 (jim_thompson5910):

that answers the second part

jimthompson5910 (jim_thompson5910):

write a rule for the nth term of this situation first term: a = 3 common ratio: r = 2 nth term \[\large a_{n} = a*r^{n-1}\] \[\large a_{n} = 3*r^{n-1}\] \[\large a_{n} = 3*2^{n-1}\]

jimthompson5910 (jim_thompson5910):

so the nth term is \[\large a_{n} = 3*2^{n-1}\]

jimthompson5910 (jim_thompson5910):

how many inches did the river rise the 7th day? plug in n = 7 and evaluate, then tell me what you get

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

that's too small

OpenStudy (anonymous):

i got 3(2) ^6

jimthompson5910 (jim_thompson5910):

that's not 36 though

jimthompson5910 (jim_thompson5910):

you evaluate 2^6 first, then you multiply

OpenStudy (anonymous):

i don't have a calculator that

jimthompson5910 (jim_thompson5910):

then use google as a calculator https://www.google.com/search?q=3%282%29+^6&ie=UTF-8

OpenStudy (anonymous):

64

jimthompson5910 (jim_thompson5910):

no it's not 64, look at the result google calculator spits out

jimthompson5910 (jim_thompson5910):

at the link i sent